

********************************************************************
UARS WEEKLY STATUS REPORT - 01 JUN 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE began a period of day only operations this week (01149) and was
subsequently turned off completely due to power concerns during the yaw
maneuver. We expect to return to normal operations shortly. Otherwise,
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM was turned off May 28; it will be turned on again June 1.
Otherwise, the instrument performed its daily solar scans for this past
week during periods of predicted UARS telemetry.
The file containing the SUSIM V19r3 MgII index, available from the SUSIM
WWW site (http://wwwsolar.nrl.navy.mil/) has been updated to include an
additional field, namely, the fraction of the day after UTC midnight
identifying the time of the MgII index measurement. This data has become
more significant recently during the present solar maximum, a time
period when the solar UV irradiance changes occur more quickly.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
>>>>> ALL COMMUNICATION TO PEM SHOULD NOW BE THROUGH THE PEM2K:: NODE <<<<<
The PEM team reports that the PEM:: node computer has failed and has
been taken off line. PEM:: appears to have hardware problems that take the
system down every few hours. Everything except the NRT automatic transfers
have been transitioned to the PEM2K:: node. The PEM:: node will be put on-line
next week for as long as possible in order to complete the NRT testing.
PEM/AXIS was activated in its full power mode on May 16 and deactivated
on May 20. At the current time, AXIS is not generating energy deposition.
SwRI is looking into this. Tests of AXIS production code on the rehost
computer are continuing.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
PEM has been working with GSFC on allowing commanding of the PEM
instrument by creating DAPs on the PEM2K machine and submitting them to the
CMS. PEM2K:: now is able to submit DAPs to the CMS.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 19 May 2001 (Orbit 52951, GMT Day 139) through 25 May 2001
(Orbit 53055, GMT Day 145).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 140/2351 52980 AXIS 1&2 OFF
ACRIM 143/2214 53024 MODE CHANGE / TEST
HRDI 144/0910 53031 Fulltime to Day-Only ops
ACRIM 144/1557 53035 Power OFF
ACRIM 144/1848 53036 Power ON
The beta angle ranged between the angles of 44.6 degrees to 21.6 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of minus 34 degrees on 14 June 2001 (DOY 165).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.2 to 1.3 degrees C. Additional battery performance data for each
battery on 18 May 2001 and 25 May 2001 are provided in Appendix D.
The clock error ranged from Ð12.0 to 12.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 23
¥ PEM = 01
¥ ACRIM = 03
¥ HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 02
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/137 show the observatory data loss to be 300 hours, 38
minutes, 10.288 seconds. This shows no change since the last report period.
This is a 0.39 percent data loss, which equals a 99.61 percent data capture for
the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The new flight software patch uplinked on 25 April 2001 continues to be
monitored this week. The FOT is pleased with its overall performance and is
fine-tuning the TAM alignment by loading new alignment matrices. This is a
slow process and it will take many alignments to lock in on the most accurate
solution.
An instrument anomaly was experienced on 05 May 2001. The instrument involved
was the ACRIM II instrument. All ACRIM telemetry for currents and voltages all
dropped to zero values suddenly. The ACRIM team in coordination with the
UARSFOT is conducting analysis and command tests to determine the cause of the
anomaly. Preliminary analysis shows that the anomaly was caused due to failed
circuitry in the ACRIM electronics box.
MEETINGS
The FOT attended a UARS Weekly Status Meeting on Monday, May 21, 2001.
The FOT attended an ISO Meeting on Monday, May 21, 2001.
The FOT attended a UARS SERB/TAC RS Meeting on Wednesday, May 23, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 30 May 2001 (DOY 150). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 25 May 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
52956 139/0822 01-315 NBTR NOOPS D-139
52962 139/1914 01-316 NBTR NOOPS D-140
52971 140/0904 01-316 NBTR NOOPS D-140
52977 140/1837 01-317 NBTR NOOPS D-141
52980 140/2351 01-318 PEM AXIS 1&2 OFF
52984 141/0640 01-317 NBTR NOOPS D-141
52992 141/1841 01-319 NBTR NOOPS D-142
52993 141/2029 01-320 OBC/TAM TAM ALIGNMENT
53000 142/0642 01-319 NBTR NOOPS D-142
53007 142/1843 01-321 NBTR NOOPS D-143
53012 143/0314 01-322 OBC Clk Adj 0.00 to 20.6 msec/day
53015 143/0644 01-321 NBTR NOOPS D-143
53022 143/1845 01-323 NBTR NOOPS D-144
53024 143/2214 01-324 ACRIM MODE CHANGE / TEST
53027 144/0216 01-325 OBC Clk Adj 20.6 to 0.00 msec/day
53030 144/0647 01-323 NBTR NOOPS D-144
53031 144/0910 01-326 HRDI Fulltime to Day-Only ops
53034 144/1417 01-327 OBC TAM alignment
53035 144/1557 01-329 ACRIM Power OFF
53036 144/1737 01-330 NBTR Clk Adj 0.00 to -20.6 msec/day
53036 144/1848 01-328 NBTR NOOPS D-145
53036 144/1848 01-331 ACRIM Power ON
53041 145/0037 01-328 NBTR NBTR PBK / RTS-58
53041 145/0037 01-333 OBC TABLE 65 (STAR CATALOG)
53042 145/0217 01-332 NBTR NOOPS D-145/NBTR TO RECORD
53043 145/0501 01-328 NBTR NOOPS D-145
53052 145/1850 01-334 NBTR NOOPS D-146
53054 145/2109 01-335 OBC Clk Adj -20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 25 May 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 25 May 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
53012 143/0317 Clk Adj 0.00 to 20.6 msec/day
53027 144/0219 Clk Adj 20.6 to 0.00 msec/day
53036 144/1742 Clk Adj 0.00 to -20.6 msec/day
53054 145/2109 Clk Adj -20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 25 May 2001
18 May 2001 - GMT Day 138 Beta = 48.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 12.0 0.000 0.0
2 +05.6/-78.4 +4.51/+3.23 +16.8/-16.0 25.4 1.030 20.5
3 +22.4/-22.4 +3.55/+2.60 +16.0/-16.0 25.4 1.020 20.5
25 May 2001 - GMT Day 145 Beta = 21.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- -------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-106.4 +5.79/+4.51 +29.6/-14.4 24.6 1.040 21.0
3 +22.4/-56.0 +4.51/+3.23 +28.8/-14.4 24.5 1.030 20.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 25 May 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-025 FOT 53049 145/1417 14170 Generic Late Acquisition
01-030 FOT 53048 145/1252 23895 Degraded Data
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-027 FOT 52928 137/1225 23880 Degraded Data
01-028 FOT 52931 137/1732 23880 Degraded Data
01-029 FOT 52947 138/1913 23882 Degraded 32K Data
01-030 FOT 53048 145/1252 23895 Degraded Data
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-025 FOT see list May 01 14170 Generic Late Acquisition
01-022 ACRIM 52709 122/2140 N/A Red Low voltage alert
01-024 ACRIM 52753 125/2045 N/A ACRIM Alerts - Power Off
SEVNTFW01/16